Use the function to evalueate the indicated expressions and simplify.
f(x) = 6X-18; f(x/3), f(x)/3

Answer by Nate(3500)

f(x) = 6x - 18
f(x/3)
f(x/3) = 6(x/3) - 18
f(x/3) = 2x - 18
f(x)/3
f(x) = (6x - 18)/3
f(x) = 2x - 6